IMPOSING A MECHANICAL MODEL ON AN IRREGULAR ELECTRONIC STATE: THE A^{1}Σ^{+} STATE OF AgH

\underline{ROBERT J. LE ROY}, DOMINIQUE R.T. APPADOO, KEVIN ANDERSON, ALIREZA SHAYESTEH, IOULI E. GORDON AND PETER F. BERNATH, Guelph-Waterloo Centre for Graduate Work in Chemistry and Biochemistry, University of Waterloo, Waterloo, Ontario N2L 3G1, Canada.

Almost since its first observation in 1925, it has been clear that the A^{1}Σ^{+} state of AgH is irregular, in that its level energy patterns and isotope effects resisted conventional systematic description.^{a} A proposal by Gerö and Schmid^{b} that this irregular behaviour was due to perturbation by the nearby B^{1}Σ^{+} state was disputed by Learner,^{c} who argued that the problem was due to an avoided crossing which gives rise to a “shelf” on the outer part of the potential well. Recent ab initio studies by Witek et al.^{d}^{,}^{e} suggest that the irregular behaviour is actually due to two avoided crossings, and that the resulting effective potential has what we call “anti-shelf” behaviour which is qualitatively different than that described by Learner.^{c} The present paper describes our efforts to determine an accurate effective potential function and delineate the degree and nature of residual non-mechanical behaviour for this state from a combined-isotopologue analysis of all available experimental data.

a E. Bengtsson and E. Olsen, Z. Physik 72, 163 (1931).

b L. Gerö and R. Schmidt, Z. Physik 121, 459 (1943).

c R.C.M. Learner, Proc. Roy. Soc. (London) A 269, 327 (1962).

d H.A. Witek, T. Nakijima and K. Hirao, J. Chem. Phys. 113, 8015 (2000).

e H.A. Witek, D.G. Fedorov, K. Hirao, A. Viel and P.-O. Widmark, J. Chem. Phys. 116, 8386 (2002).